# Dynamic Fog Computing for Enhanced LLM Execution in Medical Applications

### v1

@@ 1 Introduction
LLMs have the potential to harness the vast quantities of accessible health information, advancing healthcare objectives, reducing costs, and improving patient outcomes [4, 5]. Such models generate human-like text from provided information, effectively bridging the gap between raw data and user interpretation and transforming vast amounts of heterogeneous, structured, or unstructured data into human-legible insights. LLMs can answer questions, summarize, paraphrase, and interpret text, outperforming human experts in certain contexts [6].
@@
To address these limitations, we present a paradigm shift in the digital health approach to LLMs: a fog-computing based architecture that dynamically relocates model execution closer to user devices [9]. Our approach creates the foundation for LLM inference environments on widely available, decentralized computing assets, such as users’ mobile phones, laptops, or existing computing resources within secure, isolated hospital and clinic networks. We propose using multiple LLM execution environments to account for the capabilities of low-power resources running smaller models and the trust and financial concerns associated with remote cloud computing.
@@
To address the technical complexities associated with decentralized LLMs execution methodologies, we developed SpeziLLM, an open-source software framework containing all the necessary tools for integrating LLM-related functionality in digital health applications. We demonstrate SpeziLLM’s broad applicability across six digital health applications, showcasing its versatility in various healthcare settings.
@@ 2 Background
Patient-accessible electronic health records (PAEHRs) have been shown to improve patient outcomes, improving engagement, self-management, and informed decision-making[10, 1]. However, limited English proficiency (LEP) and health literacy disparities may hinder PAEHR utilization, even for those not facing technology or internet access barriers [11, 12]. In contrast, LLM’ sophisticated querying capabilities have the potential to enable patients to generate context-aware insights into their health data at any degree of complexity and in various languages. Preliminary validations have established the efficacy of LLMs such as OpenAI GPT models in this context [13, 14]. One such demonstration is the open-source LLMonFHIR [15] application, designed to facilitate a ”dialogue” with FHIR health information [16].
@@
Edge computing, a distributed paradigm, brings computation and data storage closer to data sources, enhancing trust, security, and privacy by processing data locally [19]. Adapting compute-intensive LLMs for edge computing can be challenging, due to the disparity in computing resources between consumer-grade mobile edge devices and specialized cloud-based infrastructures. Edge devices’ limited computational power, resulting from physical dimensions, heat dissipation, battery life, and cost constraints, restricts memory allocation and GPU capabilities. Ongoing efforts to overcome these limitations and enhance mobile inference performance include the introduction of more compact LLMs and model compression techniques (4-bit or 1-bit quantization) [20, 21, 22, 23].
@@
  • Edge Layer (Local): Low-power IoT and end-user devices at the network’s edge where data is generated and utilized.
@@ 3 Architecture
Our goal is to transparently and dynamically shift the inference environment of LLMs closer to the user’s device. We aim to provide a uniform and interchangeable interface for interacting with LLMs, regardless of execution locality.
@@
In order to establish a shared mental model, we propose a set of terminologies and concepts applicable to LLM interactions (Figure 2). This model contains four major components. A Schema serves as a blueprint for a specific model configuration, representing such parameters as model type or temperature. The schema, once initialized, is immutable, allowing for key consistency and side effect guarantees. It does not contain a variable or in-flight inference state. A Session acts as the LLM in execution, housing the current context and generation state and performing the actual inference. Sessions are instantiated via passage of the schema to a centralized Runner component, which internally delegates and assigns the session to a specific local, fog, or remote Platform. A platform is responsible for setting up the LLM execution environment and ensuring concurrency-safe access to shared resources. A session always runs on a specific platform. The runner is configured with various platforms, representing a system’s different LLM inference layers.
@@ 3.1 Retrieval-Augmented Generation Capabilities
To mitigate these challenges, major LLM service providers like OpenAI and Anthropic and openly available LLMs like Llama3 [31] have introduced function calling mechanisms [32], specific instantiations of Retrieval-Augmented Generation (RAG) [33], that enable LLMs to have structured and reliable interaction with external systems.
@@ 3.2 Dynamic LLM Task Dispatching
Figure 4 illustrates the process by which an LLM inference job is assigned to a computing resource within the fog layer. It requires two components: A fog node that advertises LLM inference services (acting as the server) and a client that discovers and consumes the LLM resource. Upon receipt of a user message, the runner discovers Fog LLM Services available within the local network. Once a service is discovered, the LLM FogSession dispatches an inference job with proper authorization credentials to the previously discovered fog LLM service via a secure connection. If the use of that resource is permitted by the Fog Auth Service, the fog node will set up the language model and stream the content back to the client, who displays the response.
Figure 4: Typical procedure LLM inference job execution within the fog layer and our established mental model as a UML sequence diagram.
@@ 4 SpeziLLM
Based on the fog computing LLM architecture described in section 3, we propose a platform enabling developer access to LLMs across edge, fog, and cloud layers. Our goal is to empower developers to integrate LLMs securely, privately, and cost-efficiently, simplifying the complexities of the decentralized architecture for digital health innovators. To that end, we introduce SpeziLLM [^1] https://github.com/StanfordSpezi/SpeziLLM/ [35]: an open-source, MIT-licensed Swift software framework offering modularized, ready-to-use LLM capabilities that can be combined strategically to suit a breadth of patient care and clinical research goals. SpeziLLM is embedded in the Stanford Spezi ecosystem [36, 37], which enables the rapid development of component-based, interoperable, and reusable digital health applications. As a first demonstration, SpeziLLM supports all major Apple operating systems, allowing seamless LLM integration across iOS, macOS, and visionOS. The framework additionally leverages Apple’s hardware and software ecosystem, integrating with HealthKit and utilizing Apple Silicon and Metal acceleration.
@@ 4.1 Cloud or remote layer
5 @Parameter(description: "Categories of health data to fetch")
@@
A full code example is available in the open-source SpeziLLM documentation [^2] https://swiftpackageindex.com/stanfordspezi/spezillm/documentation/spezillmopenai/functioncalling.
@@ 4.2 Fog layer
The client, represented by the SpeziLLM framework, initiates service discovery by advertising fog nodes (Figure 4). This resource announcement and selection process utilizes Multicast DNS (mDNS) [38] and DNS Service Discovery (DNS-SD) [39] to ensure widespread compatibility across various operating systems, including Linux through Avahi [^3] https://avahi.org/ and Apple platforms via Bonjour [^4] https://developer.apple.com/bonjour.
@@ 4.3 Edge or local layer
SpeziLLM’s local execution builds on the open-source llama.cpp library, [^7] https://github.com/ggerganov/llama.cpp which now supports all major open-weight models. The library leverages Apple’s Silicon hardware acceleration and software frameworks like Accelerate [^8] https://developer.apple.com/documentation/accelerate and Metal. [^9] https://developer.apple.com/metal/ It supports vectorization, quantization, hybrid CPU/GPU inference, and the offloading of specific inference tasks to optimized chip components. An open-source XCFramework version of llama.cpp [^10] https://github.com/StanfordBDHG/llama.cpp was forked, enabling binary distribution for all Apple platforms, proper semantic versioning, and dependency management via the Swift Package Manager (SPM).
SpeziLLM manages resource tasks for local execution, freeing developers to focus on application logic. The LLM LocalPlatform enforces one sequential execution job at a time, ensuring proper local inference. The framework also includes utility components, like an LLM download and persistence manager for efficient model file retrieval and setup.
SpeziLLM’s local execution layer supports the open-weight language models listed in Table 1 with minimal configuration. Developers must provide a LocalSchema configuration and an LLM model file in the llama.cpp GPT-Generated Unified Format (GGUF) format. [^11] https://github.com/ggerganov/ggml/blob/master/docs/gguf.md SpeziLLM’s local execution component additionally supports the integration of any GGUF format model and configuration of varying prompt structures.
@@ 4.4 Methods
We also evaluated SpeziLLM’s utility in mobile application development for those without prior experience with SpeziLLM or the Stanford Spezi ecosystem. To that end, CS students enrolled in CS342121212https://cs342.stanford.edu at Stanford University were voluntarily sampled. CS342 is a ten-week, team-based course in which students learn to design and build secure digital solutions to unmet health needs. An anonymous survey was distributed via Google Forms to 15 students whose final applications were developed using SpeziLLM. The survey consisted of eight five-point Likert-scale questions (detailed in Table 3), and was designed to assess the usability and functionality of SpeziLLM. Data collection spanned five days. Participants were notified via direct messages (with one ”reminder”) to maximize response rates. Responses were analyzed using descriptive statistics.
The primary objective of this study was to assess the usability, functionality, and adaptability of SpeziLLM among novices in Swift and mobile app development. Limitations include a small sample size, poor generalizability (the sample is composed of Computer Science students in a particular course at a particular institution), and varying levels of respondent engagement with SpeziLLM, the broader Stanford Spezi ecosystem, and the Swift language.
@@ 4.5 Results
LLMonFHIR can switch between underlying LLM inference environments, combining different LLM layers to dynamically dispatch tasks to the most suitable LLM. Simple tasks like the summarization, transformation, or interpretation of single, raw FHIR resources (6(b)) can be handled by local or fog-based models with higher trust levels (section 3), minimizing submission of potential patient identifiers to remote cloud providers. Outputs from these tasks serve as inputs for more complex tasks, such as the interactive chat view (6(c)), which builds on locally-generated summaries, allowing the cloud-based OpenAI LLM to fetch, explain, and interpret the summaries in a user dialogue.
SpeziLLM’s uniform interface and cross-layer dynamic dispatch capability are heavily utilized in the HealthGPT (3) application, rearchitected to utilize SpeziLLM and multiple LLM execution environments [44]. This open-source, Stanford Spezi-based application allows users to interactively query their Apple Health data using natural language [44] (Figure 7). Users can query metrics like sleep, step count, exercise minutes, body mass, and heart rate through a speech-to-text and text-to-speech capable chat interface. Aside from the OpenAI cloud LLM inference, HealthGPT supports models running in the local or fog layer, ensuring that sensitive health information and user conversations are processed in trusted environments.
@@
Figure 8: Comparison of inference speed and time to first token for different LLMs in HealthGPT. Each measurement represents the mean of five quantifications, with the standard deviation between these measurements in brackets. Measurements were made in response to the question ”How much did I sleep last week and how can I improve my sleep?.”
@@
SpeziLLM has been additionally instantiated in the Stanford CS342 projects Nourish (4), Stronger [45] (5), and Intake [46] (6), all of which used the framework for chat interfaces and function calling mechanisms to interface with application domain logic. The Intake project populates medical intake forms with EHR-provided FHIR record data via interactive language models. The selection and filtering of relevant FHIR records combines rules-based logic with LLM inference, addressing the limitations of rules-based logic alone in handling the complexities and customizations of the FHIR format. Additional user-friendly features include the collection of additional patient data (the primary concern/reason for visit, for example) through an interactive chat-based LLM interface.
The student survey portion of our analysis had an 86.67% response rate (n = 13). Quantitative reviews of SpeziLLM were generally positive (Figure 9). Students noted the simplicity of integrating the framework into their digital health prototypes (Q1) and praised its utility in easing the use of LLMs therein (Q5). The functionality, including the transparent execution environment and declarative OpenAI function calling DSL, received similarly high ratings (Q8). Respondents reported–albeit with the highest variation–that the learning curve was challenging (Q2). SpeziLLM’s API design, naming conventions, and documentation were thought to be intuitive and helpful (Q4, Q6). The framework’s error and exception handling capability was identified as needing improvement (Q7). Overall, satisfaction was high, with students noting that SpeziLLM significantly eased and accelerated LLM integration into their digital health applications (Q3). The time-constrained eight-week course coding period necessitated a steep learning curve, and negative feedback stemmed largely from limited software engineering experience and a lack of prior exposure to Swift, SwiftUI, and the Stanford Spezi ecosystem.
@@ 5 Discussion
Throughout the development of our case studies, we observed an inverse correlation between context size and inference speed. The memory swapping associated with larger contexts served to degrade performance, even using advanced devices like the iPhone 15 Pro with 8GB RAM  [xu2023llmcad](subsection 4.5). Similar reductions occur in fog nodes within Docker containers. The high cost of inference-capable hardware may pose a challenge to the widespread adoption of LLM capabilities.
@@
Technological advances continue to create the potential for local LLM execution on edge devices. Development of LLM inference software optimizations for mobile operating systems is ongoing, with a particular focus on improving execution in memory-limited contexts. These projects, which include Android’s AICore and Apple’s CoreML, aim to optimize large-scale model execution on standard mobile devices and with extensive contexts [21, 22, 23]. SpeziLLM’s flexible model for uniform interactions can readily incorporate advancements in local execution capability, accelerating LLM application development and deployment for digital health innovators and the patients they serve.

### v2

@@ 1 Introduction
LLMs have the potential to democratize the vast quantities of accessible health information, advancing healthcare objectives, reducing costs, and improving patient outcomes [4, 5]. Such models generate human-like text from provided information, effectively bridging the gap between raw data and user interpretation and transforming vast amounts structured or unstructured data into human-legible insights. LLMs can answer questions, summarize, paraphrase, and interpret text, outperforming human experts in certain contexts [6].
@@
To address these limitations, we present a paradigm shift in the digital health approach to LLMs: a fog-computing based architecture that dynamically relocates model execution closer to user devices [9] (see Figure 1). Our approach creates the foundation for LLM inference environments on widely available, decentralized computing assets, such as users’ mobile phones, laptops, or existing computing resources within secure, isolated hospital and clinic networks. We propose using multiple LLM execution environments to account for the limited capabilities of low-power resources running smaller models and the trust and financial concerns associated with remote cloud computing.
@@
To complement our LLM fog architecture, we additionally present SpeziLLM, an open-source software framework containing all the necessary tools for integrating LLM-related functionality in digital health applications. We demonstrate SpeziLLM’s broad applicability across six digital health applications, showcasing its versatility in various healthcare settings, and assess its usability, functionality, and adaptability via a survey of student mobile health developers. The LLM space is fast-developing, so SpeziLLM was designed to be LLM-agnostic–able to integrate with various LLMs without being tied to any specific one. As such, this paper does not evaluate the output quality of particular LLMs.
@@ 2 Background
Patient-accessible electronic health records (PAEHRs) have been shown to improve patient outcomes, improving engagement, self-management, and informed decision-making[10, 1]. However, limited English proficiency (LEP) and health literacy disparities may hinder PAEHR utilization [11, 12]. LLMs’ sophisticated querying capabilities have the potential to mitigate these barriers, enabling patients to generate context-aware insights into their health data at any degree of complexity and in various languages. Preliminary validations have established the efficacy of LLMs such as OpenAI’s GPT in this context [13, 14]. One such demonstration is the open-source LLMonFHIR application [15], designed to facilitate a ”dialogue” with FHIR health information [16].
@@
Edge computing, a distributed paradigm, brings computation and data storage closer to data sources, enhancing trust, security, and privacy by processing data locally [19]. Adapting compute-intensive LLMs for edge computing can be challenging, due to the disparity in computing resources between consumer-grade mobile edge devices and specialized cloud-based infrastructures. Edge devices’ limited computational power, resulting from physical dimensions, heat dissipation, battery life, and cost constraints, restricts memory allocation and GPU capabilities. Ongoing efforts to overcome these limitations and enhance mobile inference performance include the introduction of more compact LLMs and model compression techniques (4-bit or 1-bit quantization) [20, 21, 22, 23, 24].
@@
  • Edge Layer (Local): Low-power Internet of Things (IoT) and end-user devices at the network’s edge where data is generated and utilized.
@@ 3 Architecture
Our goal is to transparently and dynamically shift the inference environment of LLMs closer to the user’s device, depending on cardinality, trust, and computational power for each layer (Figure 1). We aim to provide a uniform and interchangeable interface for interacting with LLMs, regardless of execution locality.
@@
In order to establish a shared mental model, we propose a set of terminologies and concepts applicable to LLM interactions (Figure 2). This model contains four major components. A Schema serves as a blueprint for a specific model configuration, representing parameters such as model type or temperature. The schema, once initialized, is immutable, allowing for key consistency and side effect guarantees. It does not contain a variable or in-flight inference state. The schema is passed to a centralized Runner component, which internally delegates and assigns it to a specific local, fog, or remote Platform. The runner is configured with various platforms, representing a system’s different LLM inference layers. A platform is responsible for setting up the LLM execution environment and ensuring concurrency-safe access to shared resources. The platform turns the schema into a concrete Session that acts as the LLM in execution, housing the current context and in-flight generation state and performing the actual inference.
@@ 3.1 Retrieval-Augmented Generation Capabilities
To mitigate these challenges, major LLM service providers like OpenAI and Anthropic and openly available LLMs like Llama3 [32] have introduced function calling mechanisms [33], specific instantiations of Retrieval-Augmented Generation (RAG) [34] that enable LLMs to have structured and reliable interaction with external systems.
@@ 3.2 Dynamic LLM Task Dispatching
Figure 4 illustrates the process by which an LLM inference job is assigned to a computing resource within the fog layer. The process requires two components: A fog node that advertises LLM inference services (acting as the server) and a client that discovers and consumes the LLM resource. Upon receipt of a user message, the runner discovers Fog LLM Services available within the local network. Once a service is discovered, the LLM FogSession dispatches an inference job with proper authorization credentials to the previously discovered fog LLM service via a secure connection. If the use of that resource is permitted by the Fog Auth Service, the fog node will set up the language model and stream the content back to the client, who displays the response.
Figure 4: Typical procedure for LLM inference job execution within the fog layer and our established mental model as a UML sequence diagram.
@@ 4 SpeziLLM
Based on the fog computing LLM architecture described in section 3, we propose a platform enabling developer access to LLMs across edge, fog, and cloud layers. Our goal is to empower developers to integrate LLMs securely, privately, and cost-efficiently, simplifying the complexities of the decentralized architecture for digital health innovators. To that end, we introduce SpeziLLM [^1] https://github.com/StanfordSpezi/SpeziLLM/ [36]: an open-source, MIT-licensed Swift software framework offering modularized, ready-to-use LLM capabilities that can be combined to suit a breadth of patient care and clinical research goals. SpeziLLM is embedded in the Stanford Spezi ecosystem [37, 38], which enables the rapid development of component-based, interoperable, and reusable digital health applications. As a first demonstration, SpeziLLM supports all major Apple operating systems, allowing seamless LLM integration across iOS, macOS, and visionOS. The framework additionally leverages Apple’s hardware and software ecosystem, integrating with HealthKit and utilizing Apple Silicon and Metal [^2] https://developer.apple.com/metal/ acceleration.
@@ 4.1 Cloud or remote layer
5 @Parameter("Categories of health data")
@@
A full code example is available in the open-source SpeziLLM documentation. [^3] https://swiftpackageindex.com/stanfordspezi/spezillm/documentation/spezillmopenai/functioncalling
@@ 4.2 Fog layer
The client, represented by the SpeziLLM framework, initiates service discovery by advertising fog nodes (Figure 4). This resource announcement and selection process utilizes Multicast Domain Name Service (mDNS) [39] and DNS Service Discovery (DNS-SD) [40] to ensure widespread compatibility across various operating systems, including Linux through Avahi [^4] https://avahi.org/ and Apple platforms via Bonjour [^5] https://developer.apple.com/bonjour.
@@ 4.3 Edge or local layer
SpeziLLM’s local execution builds on the open-source llama.cpp library [^8] https://github.com/ggerganov/llama.cpp, which now supports all major open-weight models. The library leverages Apple’s Silicon hardware acceleration and software frameworks like Accelerate [^9] https://developer.apple.com/documentation/accelerate and Metal. It supports vectorization, quantization, hybrid CPU/GPU inference, and the offloading of specific inference tasks to optimized chip components. An open-source XCFramework version of llama.cpp  [^10] https://github.com/StanfordBDHG/llama.cpp was forked, enabling binary distribution for all Apple platforms, proper semantic versioning, and dependency management via the Swift Package Manager (SPM).
SpeziLLM manages resource tasks for local execution, freeing developers to focus on application logic. The LLM LocalPlatform (section 3) enforces one sequential execution job at a time, ensuring proper local inference. The framework also includes utility components, like an LLM download and persistence manager for efficient model file retrieval and setup.
SpeziLLM’s local execution layer supports the open-weight language models listed in Table 1 with minimal configuration. Developers must provide a LocalSchema configuration and an LLM model file in the llama.cpp GPT-Generated Unified Format (GGUF) format [^11] https://github.com/ggerganov/ggml/blob/master/docs/gguf.md. SpeziLLM’s local execution component additionally supports the integration of any GGUF format model and configuration of varying prompt structures.
@@ 4.4 Methods
We also evaluated SpeziLLM’s utility in mobile application development for those without prior experience with SpeziLLM or the Stanford Spezi ecosystem. To that end, CS students enrolled in CS342121212https://cs342.stanford.edu at Stanford University were voluntarily sampled. CS342 is a ten-week, team-based course in which students learn to design and build secure digital solutions to unmet health needs. An anonymous survey was distributed via Google Forms to 15 students whose final applications (Nourish (4), Stronger (5), Intake (6)) were developed using SpeziLLM in groups of five. The survey consisted of eight five-point Likert-scale questions (detailed in Table 3), and was designed to assess the usability and functionality of SpeziLLM. Data collection spanned five days. Participants were notified via direct messages (with one ”reminder”) to maximize response rates. Responses were analyzed using descriptive statistics.
The primary objective of this study was to assess the usability, functionality, and adaptability of SpeziLLM among novices in Swift and mobile app development. Limitations include a small sample size, poor generalizability (the sample is composed of Computer Science students in a particular course at a particular institution), the lack of a control group, and varying levels of respondent engagement with SpeziLLM, the broader Stanford Spezi ecosystem, and the Swift language.
@@ 4.5 Results
LLMonFHIR (Figure 6) can switch between underlying LLM inference environments, combining different LLM layers to dynamically dispatch tasks to the most suitable LLM. Simple tasks like the summarization, transformation, or interpretation of single, raw FHIR resources (6(b)) can be handled by local or fog-based models with higher trust levels (section 3), minimizing exposure of potential patient identifiers to remote cloud providers. Outputs from these tasks serve as inputs for more complex tasks, such as the interactive chat view (6(c)), which builds on locally-generated summaries, allowing the cloud-based OpenAI LLM to fetch, explain, and interpret the summaries in a user dialogue without accessing raw, potentially sensitive FHIR resources.
SpeziLLM’s uniform interface and cross-layer dynamic dispatch capability are heavily utilized in the HealthGPT (3) application (Figure 7), rearchitected to utilize SpeziLLM and multiple LLM execution environments [45]. This open-source, Stanford Spezi-based application allows users to interactively query their Apple Health data using natural language [45]. Users can query metrics like sleep, step count, exercise minutes, body mass, and heart rate through a bidirectional speech-to-text capable chat interface at any degree of complexity and in various languages (7(c)). Aside from the OpenAI cloud LLM inference (7(a)), HealthGPT supports models running in the local or fog layer (7(b)), ensuring that sensitive health information and user conversations are processed in trusted environments.
@@
Figure 8: Comparison of inference speed and time to first token for different LLMs in HealthGPT. Each measurement represents the mean of five quantifications, with the standard deviation between these measurements in brackets. Measurements were made in response to the question ”How much did I sleep last week and how can I improve my sleep?”.
@@
SpeziLLM has been additionally instantiated in the Stanford University CS342 projects Nourish (4), Stronger (5) [46], and Intake (6) [47], all of which used the framework for chat interfaces and function calling mechanisms to interface with application domain logic. The Intake project populates medical intake forms with EHR-provided FHIR record data via interactive language models. The selection and filtering of relevant FHIR records combines rules-based logic with LLM inference, addressing the limitations of rules-based logic alone in handling the complexities and customizations of the FHIR format. Additional user-friendly features include the collection of additional patient data (the primary concern/reason for visit, for example) through an interactive chat-based LLM interface.
The student survey portion of our analysis had an 86.67% response rate (n = 13) to all eight questions. Quantitative reviews of SpeziLLM were generally positive (Figure 9). Students noted the simplicity of integrating the framework into their digital health prototypes (Q1) and praised its utility in easing the use of LLMs therein (Q5). The functionality, including the transparent execution environment and declarative OpenAI function calling DSL, received similarly high ratings (Q8). Respondents reported–albeit with the highest variation–that the learning curve was challenging (Q2). SpeziLLM’s API design, naming conventions, and documentation were thought to be intuitive and helpful (Q4, Q6). The framework’s error and exception handling capability was identified as needing improvement (Q7). Overall, satisfaction was high, with students noting that SpeziLLM significantly eased and accelerated LLM integration into their digital health applications (Q3). The time-constrained eight-week course coding period necessitated a steep learning curve, and negative feedback stemmed largely from limited software engineering experience and a lack of prior exposure to Swift, SwiftUI, and the Stanford Spezi ecosystem.
@@ 5 Discussion
Throughout the development of our case studies, we observed an inverse correlation between context size and inference speed. The memory swapping associated with larger contexts served to degrade performance, even using advanced devices like the iPhone 15 Pro with 8GB RAM [49] (subsection 4.5). Similar reductions occur in fog nodes within Docker containers. The high cost of inference-capable hardware may pose a challenge to the widespread adoption of LLM capabilities.
@@
Technological advances continue to create the potential for local LLM execution on edge devices. Development of LLM inference software optimizations for mobile operating systems is ongoing, with a particular focus on improving execution in memory-limited contexts. These projects, which include Android’s AICore and Apple’s CoreML / Intelligence, aim to optimize large-scale model execution on standard mobile devices and with extensive contexts [21, 22, 23, 24]. SpeziLLM’s flexible model for uniform interactions can readily incorporate advancements in local execution capability, accelerating LLM application development and deployment for digital health innovators and the patients they serve.